OpenStudy (anonymous):

Discuss the symmetry of the graph of the polynomial function. f(x)=x^5-x

OpenStudy (ranga):

This is an odd function. How? f(x) = x^5 - x f(-x) = (-x)^5 - (-x) = -x^5 + x = -(x^5 - x) = -f(x) So whenever f(-x) = -f(x) it is an odd function. Odd functions are symmetric about the origin.

OpenStudy (anonymous):

I see-thx

OpenStudy (ranga):

You are welcome. FYI: Even function is when f(-x) = f(x). Example f(x) = x^2. Even functions are symmetric about the y-axis.

OpenStudy (anonymous):

Very helpful, thx